OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $78 for 1 adult ticket and 2 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 78 = x + (x - 6) What is the regular price of a child's ticket? $50 $56 $75 $81

OpenStudy (anonymous):

I think there is something wrong with your equation. 78dollars is the price of 1 adult ticket and 2 children ticket. So, 78 = x + 2(x - 6) since 78 is the amount cost during off peak, during regular periods, it will cost 78*2= 156dollars 156=x+2(x-6) 156= 3x-12 3x= 168 x=56 Child ticket cost, 56-6=50 dollars

OpenStudy (anonymous):

Thanks!!